# Commutation of Smyth and Hoare Power Constructions in Well-filtered Dcpos

### v3

@@ Abstract
t1t1footnotetext: This work is supported by the National Natural Science Foundation of China (No.12231007)
@@ 1 Introduction
Non-deterministic choices are important semantic concepts that offer new insights in designing more powerful programming languages. They were originally modeled by the power constructions on the category of domains. The most common of these are the Hoare and the Smyth powerdomains [22], which can respectively serve as the denotational semantics for angelic and demonic non-deterministic choices. Notably, the Hoare and the Smyth powerdomains also give rise to monads. In 1991, Eugenio Moggi showed that, within the category of sets, the collection of all finite sets could serve as denotations for nondeterministic programming languages [19]. He further proved that this structure forms a monad on the category of sets, thereby linking power constructions with monads. Over time, monads have evolved into a fundamental tool for constructing denotational semantic models in functional programming languages.
@@
In 1969, Jon Beck proposed the notion of the distributive law between two monads defined on the same category [2]. Such a distributive law ensures that the composition of these two monads remains a monad, providing a new approach to establishing denotational semantic models in functional programming languages. Moreover, the two composite monads obtained under different composition orders may differ; if they coincide, i.e., the monad functors induce isomorphic mappings on objects, then the monads are said to obey a commutative law. Such commutativity ensures that the order of program execution becomes immaterial in computation. This practical significance raises two fundamental problems: (1) Across which categories does a distributive law exist between the two denotational semantic models for nondeterministic programming languages the Hoare power monad and the Smyth power monad? (2) Within a specified category, under what conditions do these two monads commute?
@@
In domain theory, numerous scholars have conducted in-depth investigations into the above two problems. In 1990, Kevin Flannery and Johannes Martin employed information systems to demonstrate the commutativity between the Hoare power monad and the Smyth power monad in the category of bounded complete algebraic domains [6]. Subsequently, Reinhold Heckmann used an entirely distinct method to prove that for any full subcategory of the category of dcpos, the commutativity of Hoare and Smyth power monads is guaranteed for variants of the two monads that are defined as the free inflationary semilattice and the free deflationary semilattice, respectively [12]. Within the category of domains, the Hoare powerdomain induces the free upper semilattices, while the Smyth powerdomain yields the free meet-continuous semilattices [9]. Here, the upper semilattices are actually inflationary semilattices, and the meet-continuous semilattices are deflationary semilattices. Thus the Hoare power monad and the Smyth power monad commute in the category of domains.
@@
Note that many categories have the category of domains as a full subcategory. The versions of Hoare and Smyth power monads considered in [12], i.e. defined as free constructions, are well-defined for all dcpos and do commute on the category of all dcpos. However, the specific free constructions here are not clear. So the versions of the power monads defined via closed and saturated compact subsets are different from the free constructions in general. It implies that the restriction to well-filtered dcpos (called $\mathcal{U}_{K}$-admitting in [11]) is necessary to obtain a well-behaved Smyth contruction. This issue has been evolving continuously. Concretely, Heckmann in [11] investigated the commutativity between the Hoare power construction $\mathcal{L}$ and a modified Smyth power construction $\mathcal{U}_{S}$ on the category of $\mathcal{U}_{S}$-admitting dcpos [11]. In this work, the extended Smyth power construction is determined by all strongly compact subsets, not the traditional Smyth power construction based on compact subsets. Thus, the $\mathcal{U}_{S}$-admissibility is strictly weaker than well-filteredness. Heckmann introduced the property of being $\mathcal{U}_{S}$-conformal and showed that for a dcpo $L$, $L$ is $\mathcal{U}_{S}$-conformal if and only if $\mathcal{L}(\mathcal{U}_{S}(L))$ is isomorphic to $\mathcal{U}_{S}(\mathcal{L}(L))$, i.e., established the sufficient and necessary conditions for the commutativity. Meanwhile, they also pointed out that the corresponding commutativity could not be characterized in the category of well-filtered dcpos.
@@
In this paper, we focus on the category of well-filtered dcpos and investigate the commutativity between the associated Hoare and Smyth power monads. In Section 3, we define the Hoare and Smyth constructions for a well-filtered dcpo $L$ as follows: The Hoare construction $\mathcal{H}$ is the set of all Scott closed subsets of $L$, ordered by inclusion; the Smyth construction $\mathcal{Q}$ is the set of all Scott compact saturated subsets of $L$, ordered by reverse inclusion order. We explicitly show that $\mathcal{Q}$ does not yield a free construction for the category $\mathbf{WF}$ of well-filtered dcpos with Scott-continuous maps, which confirms the significance and necessity of our research. In Section 4, we introduce a property called $(\mathrm{KC})$, which is weaker than the consonance in terms of posets. Based on this we proved that $\mathcal{H}$ and $\mathcal{Q}$ commute in the sense of that $\mathcal{HQ}(L)$ is isomorphic to $\mathcal{QH}(L)$ under a certain function, if and only if $L$ satisfies the property $(\mathrm{KC})$ and the Scott topology coincides with the upper Vietoris topology on $\mathcal{Q}(L)$. It is necessary to point out that in [4], de Brecht and Kawai also provided a Smyth construction via compact saturated subsets and investigated the commutativity between the Smyth construction and the Hoare construction. However, our work differs from theirs, since they consider the category of quasi-polish topological spaces while we explore that of well-filtered dcpos.
@@
Notably, unlike Heckmann’s modified Smyth power construction $\mathcal{U}_{S}$, $\mathcal{Q}$ in this paper induces the Smyth power monad on $\mathbf{WF}$. This relies crucially on Xu’s result [26] that the Smyth power space endowed with the Scott topology of any well-filtered space remains well-filtered. Furthermore, since every complete lattice with the Scott topology is well-filtered [25, Corollary 3.2], we conclude that $\mathcal{H}$ also forms a monad on $\mathbf{WF}$. Thus in Section 5, we analyze the properties of the composite monad. The morphism $\phi$ defined in Definition 4.1 is shown to be a distributive law, which ensures that the composition $\mathcal{QH}$ remains a monad on $\mathbf{WF}$. Finally, motivated by the growing interest in Eilenberg-Moore algebras, we characterize the Eilenberg-Moore category of $\mathcal{QH}$ as a subcategory of Frm the category of frames and frame homomorphisms.
@@ 2 Preliminaries
In this section, we will give a brief introduction to some basic concepts and notions to be used in this paper. One can refer to [1, 9, 10] for more details.
Given a poset $P$, $A\subseteq P$ is an upper set if $A={\uparrow}A$, where ${\uparrow}A=\{x\in P:x\geq a\ \mathrm{for\ some}\ a\in A\}$. The lower set is defined dually. The subset $D$ of $P$ is directed if every finite subset of $D$ has an upper bound in $D$. If each directed subset in a poset $P$ has a supremum, we call $P$ is a directed complete partially ordered set (dcpo). A subset $U$ of $P$ is called Scott open iff it is an upper set and for each directed subset $D$ whose supremum exists, written as $\sup D$, $\sup D\in U$ implies $D\cap U\neq\emptyset$. The complement of a Scott open set is called Scott closed. The collection of all Scott open subsets, denoted by $\sigma(P)$, forms a topology called Scott topology and in general, we write $\Sigma P=(P,\sigma(P))$. Besides, we use $\Gamma(P)$ to denote the set of all Scott closed subsets of $P$. A continuous map between two posets equipped with Scott topology is called Scott-continuous. Generally, $\mathbf{DCPO}$ is used to denote the category of all dcpos with Scott-continuous mappings.
Let $X$ be a topological space. We use $\mathcal{O}(X)$ denote the collection consisting of all open subsets of $X$. When $X$ is a $T_{0}$ space, the specialization preorder $\leq$ defined by $x\leq y$ iff $x\in cl(\{y\})$ is a partial order and is called specialization order. A subset of $X$ is saturated if it is the intersection of open sets, or equivalently, it is an upper set in the order of specialization. $X$ is called well-filtered if for each filter basis $\mathcal{C}$ of compact saturated subsets and any open subset $U\in\mathcal{O}(X)$ with $\bigcap\mathcal{C}\subseteq U$, $K\subseteq U$ for some $K\in\mathcal{C}$. We usually say a dcpo $P$ is well-filtered if it equipped with the Scott topology is a well-filtered space. Furthermore, a space $X$ is called coherent if the intersection of any two compact saturated sets of $X$ is again compact.
In this paper, we denote the set of all compact saturated subsets (including the empty set) and that of all closed subsets of a space $X$ by $\mathcal{Q}(X)$ and $\Gamma(X)$, respectively. The topology on $\mathcal{Q}(X)$ generated by $\{\square U:U\in\mathcal{O}(X)\}$ as a base, where $\Box U=\{K\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{Q}(X)}:K\subseteq U\}$, is called the upper Vietoris topology and denoted by $\upsilon({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{Q}(X)})$. The upper topology on $\Gamma(X)$ coincides with the lower Vietoris topology, which is written as $\nu(\Gamma(X))$ and generated by $\{\lozenge U:U\in\mathcal{O}(X)\}$ as a subbase, where $\lozenge U=\{A\in\Gamma(X):A\cap U\neq\emptyset\}$.
@@
$\mathcal{T}$$\mathcal{T}^{2}$$\mathcal{T}$$\mathcal{T}$$\eta_{\mathcal{T}}$$\mathcal{T}\eta$$\mu$$\mathcal{T}^{3}$$\mathcal{T}^{2}$$\mathcal{T}^{2}$$\mathcal{T}$$\mathcal{T}\mu$$\mu\mathcal{T}$$\mu$$\mu$
@@
The category composed of $\mathcal{T}$-algebras and their homomorphisms is called Eilenberg-Moore category of $\mathcal{T}$ over $\mathbf{C}$.
@@
    If $(C,\alpha)$ is a $\mathcal{T}$-algebra, then $\alpha:(\mathcal{T}C,\mu_{C})\rightarrow(C,\alpha)$ is a $\mathcal{T}$-homomorphism.
@@
    If $h_{1},h_{2}:(\mathcal{T}C,\mu_{C})\rightarrow(C^{\prime},\alpha)$ are $T$-homomorphisms with $h_{1}\circ\eta_{C}=h_{2}\circ\eta_{C}$, then $h_{1}=h_{2}$.
@@ 3.1 The Hoare power construction
For any collection $\{A_{i}:i\in I\}\subseteq\mathcal{H}(L)$, since $\mathcal{H}(f)$ clearly is order preserving, $\sup_{i\in I}\mathcal{H}(f)(A_{i})\subseteq\mathcal{H}(f)(\sup_{i\in I}A_{i})$ holds. Meanwhile,
@@
where the second symbol $\subseteq$" is guaranteed by the continuity of $f$. Thus we conclude that $\mathcal{H}(f)$ preserves arbitrary suprema. Additionally, one can verify that $\mathcal{H}$ preserves the identities and the compositions of any two morphisms. Along with the above, we can obtain the following result.
@@
It follows from [25, Corollary 3.2] that each complete lattice endowed with the Scott topology is well-filtered. So there is a forgetful functor from $\mathbf{CL}$ to $\mathbf{WF}$. Andrea Schalk has proven that the functor $P_{D}^{b}:\mathbf{DCPO}\rightarrow\mathbf{CL}$ is left adjoint to the forgetful functor $U:\mathbf{CL}\rightarrow\mathbf{DCPO}$ (see [21, Theorem 6.2]). Note that if the functor $P_{D}^{b}$ is restricted from the category $\mathbf{DCPO}$ to its full subcategory $\mathbf{WF}$, it is just the functor $\mathcal{H}$. This immediately yields the following result.
@@
Since each pair of adjunction determines a monad, [1, Proposition 4.2.1], we deduce that the triple $(U\circ\mathcal{H},\eta,U\varepsilon_{\mathcal{H}})$, where $\eta$ and $\varepsilon$ are the unit and counit respectively, turns into a monad on $\mathbf{WF}$. For ease of expressions, we will abbreviate $U\circ\mathcal{H}$ as $\mathcal{H}$. Meanwhile, when $U\varepsilon_{\mathcal{H}}$ works on a well-filtered $\mathrm{dcpo}$ $L$, one can calculate $U\varepsilon_{\mathcal{H}(L)}$ as the natural transformation from $\mathcal{H}(\mathcal{H}(L))$ to $\mathcal{H}(L)$ that maps $\mathcal{A}$ to $\sup_{\mathcal{H}(L)}\mathcal{A}$.
@@
It follows from this lemma that $\sup_{\mathcal{H}(L)}\mathcal{A}=\bigcup\mathcal{A}$. Now we conclude that
@@
The endofunctor $\mathcal{H}$ together with the unit $\eta:Id\rightarrow U\circ\mathcal{H}$ and the multiplication $\mu:\mathcal{H}^{2}\rightarrow\mathcal{H}$ forms a monad on $\mathbf{WF}$, where $\eta$ and $\mu$ are defined concretely for a well-filtered dcpo $L$ as follows:
$\eta_{L}(x)={\downarrow}x$, for any $x\in L$
@@
Combining the fact that $\mathbf{CL}$ is the category of unital inflationary semilattices for $\mathbf{WF}$ with that the Hoare power well-filtered dcpo has the universal property, we have that
@@
Based on their definitions, it is easy to verify them.

### v5

@@ Abstract
t1t1footnotetext: This work is supported by the National Natural Science Foundation of China (No. 12231007), the National Natural Science Foundation of China (No. 5kj20250124) and the Natural Science Foundation of Shandong Province, China (No. ZR2024QA195).
@@ 1 Introduction
Non-deterministic choices are important semantic concepts that offer new insights in designing more powerful programming languages. They were originally modeled by the power constructions on the category of domains. The most common of these are the Hoare and the Smyth powerdomains [22], which can respectively serve as the denotational semantics for angelic and demonic non-deterministic choices. Notably, the Hoare and the Smyth powerdomains also give rise to monads. In 1991, E. Moggi showed that, within the category of sets, the collection of all finite sets could serve as denotations for nondeterministic programming languages [19]. He further proved that this structure forms a monad on the category of sets, thereby linking power constructions with monads. Over time, monads have evolved into a fundamental tool for constructing denotational semantic models in functional programming languages.
@@
In 1969, J. Beck proposed the notion of the distributive law between two monads defined on the same category [2]. Such a distributive law ensures that the composition of these two monads remains a monad, providing a new approach to establishing denotational semantic models for functional programming languages. Moreover, the two composite monads obtained under different composition orders may differ; if they coincide, i.e., the monad functors induce isomorphic mappings on objects, then the monads are said to obey a commutative law. Such commutativity ensures that the order of program execution becomes immaterial in computation. This practical significance raises two fundamental problems: (1) Across which categories does a distributive law exist between the two denotational semantic models for nondeterministic programming languages – the Hoare power monad and the Smyth power monad? (2) Within a specified category, under what conditions do these two monads commute?
@@
In domain theory, numerous scholars have conducted in-depth investigations into the above two problems. In 1990, K. Flannery and J. Martin employed information systems to demonstrate the commutativity between the Hoare power monad and the Smyth power monad in the category of bounded complete algebraic domains [6]. Subsequently, R. Heckmann used an entirely distinct method to prove that for any full subcategory of the category of dcpos, the commutativity of Hoare and Smyth power monads is guaranteed for variants of the two monads that are defined as the free inflationary semilattice and the free deflationary semilattice, respectively [12]. Within the category of domains, the Hoare powerdomain induces the free up-complete $\vee$-semilattices, while the Smyth powerdomain yields the free meet-continuous semilattices [9]. Here, the up-complete $\vee$-semilattices are actually inflationary semilattices, and the meet-continuous semilattices are deflationary semilattices. Thus the Hoare power monad and the Smyth power monad commute in the category of domains.
@@
Note that many categories have the category of domains as a full subcategory. The versions of Hoare and Smyth power monads considered in [12], i.e. defined as free constructions, are well-defined for all dcpos and do commute on the category of all dcpos. As shown by A. Schalk in [21], we know the Hoare power construction defined by the Scott closed subsets does yield the free unital inflationary semilattice in $\mathbf{DCPO}$; however, it is not clear what constructions can give a free deflationary semilattice in $\mathbf{DCPO}$. We know that the well-filteredness guarantees that the collection of all compact saturated subsets with the reverse inclusion order is a dcpo [9]. So the issue of restricting to well-filtered dcpos (called $\mathcal{U}_{K}$-admitting dcpos in [11]) in order to obtain a well-behaved Smyth construction has attracted increasing attention. Concretely, R. Heckmann in [11] investigated the commutativity between the Hoare power construction $\mathcal{L}$ and a modified Smyth power construction $\mathcal{U}_{S}$ on the category of $\mathcal{U}_{S}$-admitting dcpos [11]. In this work, the modified Smyth power construction is determined by all strongly compact subsets, which differs from the traditional Smyth power construction based on compact subsets. Thus, the $\mathcal{U}_{S}$-admissibility is strictly weaker than well-filteredness. R. Heckmann introduced the property of being $\mathcal{U}_{S}$-conformal and showed that for a dcpo $L$, $L$ is $\mathcal{U}_{S}$-conformal if and only if $\mathcal{L}(\mathcal{U}_{S}(L))$ is isomorphic to $\mathcal{U}_{S}(\mathcal{L}(L))$, i.e., established the sufficient and necessary conditions for the commutativity. Meanwhile, they also pointed out that they could not show an analogous commutativity result for $\mathcal{L}$ and the Smyth power construction based on compact subsets.
@@
In this paper, we focus on the category of well-filtered dcpos and investigate the commutativity between the associated Hoare and Smyth power monads. In Section 3, we define the Hoare and Smyth constructions for a well-filtered dcpo $L$ as follows: The Hoare construction $\mathcal{H}$ is the set of all Scott closed subsets of $L$, ordered by inclusion; the Smyth construction $\mathcal{Q}$ is the set of all Scott compact saturated subsets of $L$, ordered by reverse inclusion order. We explicitly show that $\mathcal{Q}$ does not yield a free construction for the category $\mathbf{WF}$ of well-filtered dcpos with Scott-continuous maps, which confirms the significance and necessity of our research. In Section 4, we introduce a property called $(\mathrm{KC})$, which is weaker than the consonance in terms of posets. Based on this, we proved that $\mathcal{H}$ and $\mathcal{Q}$ commute in the sense of that $\mathcal{HQ}(L)$ is isomorphic to $\mathcal{QH}(L)$ under a certain function, if and only if $L$ satisfies the property $(\mathrm{KC})$ and the Scott topology coincides with the upper Vietoris topology on $\mathcal{Q}(L)$. It is necessary to point out that in [4], M. de Brecht and T. Kawai also provided a Smyth construction via compact saturated subsets and investigated the commutativity between the Smyth construction and the Hoare construction. However, our work differs from theirs, since they consider the category of topological spaces while we explore that of well-filtered dcpos.
@@
Notably, unlike R. Heckmann’s modified Smyth power construction $\mathcal{U}_{S}$, $\mathcal{Q}$ induces the Smyth power monad on $\mathbf{WF}$. This relies crucially on Xu’s result [26] that the Smyth power space endowed with the Scott topology of any well-filtered space remains well-filtered. Meanwhile, since every complete lattice with the Scott topology is well-filtered [25, Corollary 3.2], we conclude that $\mathcal{H}$ also gives rise to a monad on $\mathbf{WF}$. Consequently, in Section 5, we show that there is a distributive law between the two monads $\mathcal{Q}$ and $\mathcal{H}$, which ensures that the composition $\mathcal{QH}$ remains a monad on $\mathbf{WF}$. Finally, motivated by the growing interest in Eilenberg-Moore algebras, we characterize the Eilenberg-Moore category of $\mathcal{QH}$ as a subcategory of Frm – the category of frames and frame homomorphisms.
@@ 2 Preliminaries
In this section, we briefly introduce some basic concepts and notations that will be used in this paper. One can refer to [1, 9, 10] for more details.
Given a poset $P$, $A\subseteq P$ is an upper set if $A={\uparrow}A$, where ${\uparrow}A=\{x\in P:x\geq a\ \mathrm{for\ some}\ a\in A\}$. The lower set is defined dually. A subset $D$ of $P$ is directed if it is nonempty and every finite subset of $D$ has an upper bound in $D$. If each directed subset in $P$ has a supremum, we call $P$ a directed complete partially ordered set (dcpo). A subset $U$ of $P$ is called Scott open iff it is an upper set and for each directed subset $D$ whose supremum exists, written as $\sup D$, $\sup D\in U$ implies $D\cap U\neq\emptyset$. The complement of a Scott open set is called Scott closed. The collection of all Scott open subsets, denoted by $\sigma(P)$, forms a topology called Scott topology and in general, we write $\Sigma P=(P,\sigma(P))$. Besides, we use $\Gamma(P)$ to denote the set of all Scott closed subsets of $P$. A continuous map between two posets equipped with Scott topology is called Scott-continuous. Generally, $\mathbf{DCPO}$ is used to denote the category of all dcpos with Scott-continuous mappings.
Let $X$ be a topological space. We write $\mathcal{O}(X)$ for the collection of open subsets of $X$. When $X$ is a $T_{0}$ space, the specialization preorder $\leq$ defined by $x\leq y$ iff $x\in cl(\{y\})$ is a partial order and is called specialization order. A subset of $X$ is saturated if it is the intersection of open sets, or equivalently, it is an upper set in the order of specialization. $X$ is called well-filtered if for each filter basis $\mathcal{C}$ of compact saturated subsets and any open subset $U\in\mathcal{O}(X)$ with $\bigcap\mathcal{C}\subseteq U$, $K\subseteq U$ for some $K\in\mathcal{C}$. We usually say a dcpo $P$ is well-filtered if it, equipped with the Scott topology, is a well-filtered space. Furthermore, a space $X$ is called coherent if the intersection of any two compact saturated subsets of $X$ is also compact.
In this paper, we denote the set of all compact saturated subsets (including the empty set) and that of all closed subsets of a space $X$ by $\mathcal{Q}(X)$ and $\Gamma(X)$, respectively. The topology on $\mathcal{Q}(X)$ generated by $\{\square U:U\in\mathcal{O}(X)\}$ as a base, where $\Box U=\{K\in\mathcal{Q}(X):K\subseteq U\}$, is called the upper Vietoris topology and denoted by $\upsilon(\mathcal{Q}(X))$. The upper topology on $\Gamma(X)$ coincides with the lower Vietoris topology, which is written as $\nu(\Gamma(X))$ and generated by $\{\lozenge U:U\in\mathcal{O}(X)\}$ as a subbase, where $\lozenge U=\{A\in\Gamma(X):A\cap U\neq\emptyset\}$.
Regarding the inflationary (deflationary) semilattices, there are a formal definition given in [9], and a concrete description provided by A. Schalk in [21]. Combining these two forms, we give the following definition:
@@
$\mathcal{T}$$\mathcal{T}^{2}$$\mathcal{T}$$\mathcal{T}$$\eta_{\mathcal{T}}$$\mathcal{T}\eta$$\mu$$\mathcal{T}^{3}$$\mathcal{T}^{2}$$\mathcal{T}^{2}$$\mathcal{T}$$\mathcal{T}\mu$$\mu_{\mathcal{T}}$$\mu$$\mu$
@@
The category composed of $\mathcal{T}$-algebras and $\mathcal{T}$-algebra homomorphisms is called Eilenberg-Moore category of $\mathcal{T}$ over $\mathbf{C}$.
@@
    If $(C,\alpha)$ is a $\mathcal{T}$-algebra, then $\alpha:(\mathcal{T}C,\mu_{C})\rightarrow(C,\alpha)$ is a $\mathcal{T}$-algebra homomorphism.
@@
    If $h_{1},h_{2}:(\mathcal{T}C,\mu_{C})\rightarrow(C^{\prime},\alpha)$ are $\mathcal{T}$-algebra homomorphisms with $h_{1}\circ\eta_{C}=h_{2}\circ\eta_{C}$, then $h_{1}=h_{2}$.
@@ 3.1 The Hoare power construction
Let $\{A_{i}:i\in I\}$ be an arbitrary subset of $\mathcal{H}(L)$. Since $\mathcal{H}(f)$ is clearly order-preserving, $\sup_{i\in I}\mathcal{H}(f)(A_{i})\subseteq\mathcal{H}(f)(\sup_{i\in I}A_{i})$ holds. Meanwhile,
@@
where the symbol $\subseteq$" is guaranteed by the continuity of $f$. Thus we conclude that $\mathcal{H}(f)$ preserves arbitrary suprema. Additionally, one can verify that $\mathcal{H}$ preserves the identities and the compositions of any two morphisms. Along with the above, we can obtain the following result.
@@
It follows from [25, Corollary 3.2] that each complete lattice endowed with the Scott topology is well-filtered. So there is a forgetful functor from $\mathbf{CL}$ to $\mathbf{WF}$. A. Schalk has proved that the functor $P_{D}^{b}:\mathbf{DCPO}\rightarrow\mathbf{CL}$ is left adjoint to the forgetful functor $U:\mathbf{CL}\rightarrow\mathbf{DCPO}$ (see [21, Theorem 6.2]). Note that if the functor $P_{D}^{b}$ is restricted from the category $\mathbf{DCPO}$ to its full subcategory $\mathbf{WF}$, it is just the functor $\mathcal{H}$. This immediately yields the following result.
@@
Since each pair of adjunctions determines a monad, [1, Proposition 4.2.1], we deduce that the triple $(U\circ\mathcal{H},\eta,U\varepsilon_{\mathcal{H}})$, where $\eta$ and $\varepsilon$ are the unit and counit respectively, turns into a monad on $\mathbf{WF}$. For ease of expressions, we will abbreviate $U\circ\mathcal{H}$ as $\mathcal{H}$. Meanwhile, when $U\varepsilon_{\mathcal{H}}$ acts on a well-filtered $\mathrm{dcpo}$ $L$, one can calculate $U\varepsilon_{\mathcal{H}}(L)$ as the natural transformation $\mu_{L}$ from $\mathcal{H}(\mathcal{H}(L))$ to $\mathcal{H}(L)$ that maps $\mathcal{A}$ to $\sup_{\mathcal{H}(L)}\mathcal{A}$.
@@
It follows from this lemma that $\sup_{\mathcal{H}(L)}\mathcal{A}=\bigcup\mathcal{A}$. So we conclude that
@@
The endofunctor $\mathcal{H}$ together with the unit $\eta:Id\rightarrow\mathcal{H}$ and the multiplication $\mu:\mathcal{H}^{2}\rightarrow\mathcal{H}$ forms a monad on $\mathbf{WF}$, where $\eta$ and $\mu$ are defined concretely for a well-filtered dcpo $L$ as follows:
$\eta_{L}(x)={\downarrow}x$, for any $x\in L$,
@@
Since $\mathbf{CL}$ is the category of unital inflationary semilattices for $\mathbf{WF}$, and Proposition 3.3 shows that the Hoare power construction $\mathcal{H}(L)$ has the universal property for every well-filtered dcpo $L$, we obtain the following result.
@@
Based on their definitions, it is easy to verify the claims.